If 493.45 grams of CO2 are produced, how many grams of C2H2 were reacted?


1
Expert's answer
2021-01-21T03:49:15-0500

C2H2 +2.5O2 = 2CO2 + H2O

n(CO2)= 493.45/44=11.215 mole

n(C2H2)=11.215/2=5.61 mole

m(C2H2)=5.61*(12+12+1+1)=145.86 grams

Answer: 145.86 grams of C2H2 reacted.